OpenStudy (anonymous):

The first term of an arithmetic series is 5 and the 22nd term is 68. Find the sum of the series. i have no clue how to do this and it has been bugging me

OpenStudy (anonymous):

here first term=5 and the last term i.e 22nd term is68 here last term=22 so by formulae sum of arithmatic series=n/2*(a+l) =22/2*(5+68) =11*73=803

OpenStudy (anonymous):

thank you so much!